OpenStudy (anonymous):

If P is the probability of a certain event occurring, which of the following describes the range of numbers P can be?

OpenStudy (anonymous):

a.\[0 \le P \le 100\] b.\[0 < P < 100\] c.\[0 \le P < 1\] d.\[0 \le P \le 1\]

Directrix (directrix):

Go here: http://www.regentsprep.org/regents/math/algebra/apr5/LProb.htm and read the information in the black/yellow chart. Then, post what you think the answer is, okay? @pedro77

OpenStudy (anonymous):

A?

Directrix (directrix):

That is correct. A probability of zero indicates that an event will not happen at all. A probability of 1 indicates that the event is a sure thing; it will happen.